OpenStudy (anonymous):

writea polynomial equation with x=2/3 and x=-5 as integer coefficient

OpenStudy (anonymous):

I'm not sure I understand the question. Does it mean to write a polynomial with integer coefficients that has x=2/3 and x=-5 as solutions?

OpenStudy (anonymous):

If so, simplify (x-2/3)(x+5)=0 and multiply by the lowest common denominator to clear the fractions.